Operations Report – May 2nd

Crew 315 Operations Report 02-05-2025

SOL: 12

Name of person filing report: Michael Andrews

Non-nominal systems: Catalytic Heater

Notes on non-nominal systems: I have officially given up on the catalytic heater. I want to make sure I don’t introduce unnecessary gas into the station or accidentally damage the heater. I am deferring repair and troubleshooting to Mission Support.

ROVERS

Spirit rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Opportunity rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Curiosity rover used: Yes

Hours: 318.8 hours (0.3 hours used)

Beginning charge: 100%

Ending charge: 89%

Currently charging: No

Perseverance rover used: Yes

Hours: 357.3 hours (0.5 hours used)

Beginning charge: 100%

Ending charge: 81%

Currently charging: No

General notes on rovers: None

Summary of Hab operations: The team had a light breakfast, began to clean the station, welcomed CNBC for a tour and individual interviews, and performed one EVA with the CNBC cameraman.

Water Use

Time of measurements: 1615
1) Per formula: 22.87 gallons
2) Smart Home Dashboard: 29.5 gallons

Water (static tank, remaining gallons): 148.10 gallons

Static tank pipe heater: Off

Static tank heater: Off

Toilet tank emptied: No

Summary of internet: Used to complete reports, learn languages, share team photos via Google drive, researching projects, and complete professional work.

Summary of suits and radios:

· EVA 21: Suits 7, 8, 9, 11 and Radios 1 ,2, 3, 4

Summary of GreenHab operations: Started to clean up the GreenHab for closeout operations, watered plants and fungal project.

WATER USE: 3.15 gallons

Heater: Off

Supplemental light: Off

Harvest: Preparing the Mission Support plants-to-go

Summary of ScienceDome operations: Started to clean up the Dome and prepare for closeout.

Dual split: Off

Summary of RAM operations: Started to clean up the RAM and prepare for closeout. All outstanding tools from previous projects and EVAs were returned to RAM inventory.

Summary of any observatory issues: None

Summary of health and safety issues: None

Questions, concerns and requests to Mission Support: Thanks for a great time!

GreenHab Report – May 2nd

Crew 315 GreenHab Report 02-05-2025

GreenHab Officer: Elena Saavedra Buckley

Environmental control (fan & heater): On

Average temperatures (last 24h): 98

Maximum temperature (last 24h): 124 (!)

Minimum temperature (last 24h): 72

Hours of supplemental light: Disabled

Daily water usage for crops: 0 gal

Daily water usage for research and/or other purposes: 0 gal

Water in Blue Tank (200 gallon capacity): N/A (drained)

Time(s) of watering for crops: N/A

Changes to crops: Gone!

Narrative: We saved two tomato plants and some green onions, parsley, and basil for Mission Support. Hoping they survive! It’s hot out. The max temperature was because we had to turn off the fan for a CNBC interview, and it took a while for the automated service to come back on.

Harvest: N/A

Support/supplies needed: None

Journalist Report – May 1st

Journalist report sol 11

At 10:30 pm last night, our HSO heard the clang of our mousetrap and sprung into action with blue painter’s tape. The mouse was still in his cell by morning, with the bait of funfetti cake crumbs turned into dust by his reasonable anxiety. Before we could take him outside, we had a chatty breakfast while discussing tomorrow’s visit from CNBC, as well as technical things like mission summaries, GreenHab teardowns, and Hab clean-ups.

At 9, the Millennial EVA Crew (Engineer, HSO, Journalist) went to Candor Chasma and began the EVA by releasing the mouse. He had become strangely wet in his cage. He fell, cowering, out of the trap and darted behind a bush, and we saw him bound up the nearest tall hill and crest over its edge with Looney Toon speed. We wished him well and hoped that last night’s rains would give him a head start, and then we descended into the riverbed and hiked through the soft sandy floor. The HSO hoped to use her drone to weave in and out of the canyon, but after some attempts it seemed that the dust and many flights had gotten to it. (Some canned air back at the Hab cleared it up.) Upon our return we realized that no more EVAs would occur without a slight sim-breaking presence. While we were gone, our Crew Artist started tearing down the GreenHab, and our Commander spoke to the CNBC journalist who will be interviewing and hosting the segment in which we’ll participate.

The afternoon was quiet. We finished our final 100cameras module and watched 2010: The Year We Make Contact, which clarified many storylines for some and seemed wildly ‘80s to all. Dinner was quick, the dessert rice pudding was coconut, and our sense of the mission suddenly sprinting toward the finish line. Make it stop!

Sol Summary – May 1st

Crew 315 Sol Summary Report 01-05-2025

Sol: 11

Summary Title: Martian Sol Mates

Author’s name: Michael Andrews

Mission Status: Nominal

Sol Activity Summary:

Crew 315 has just completed their penultimate day on Mars. And for the mouse that has plagued our kitchen over the last 2 sols, it was his last sol on Mars. He was captured in our mouse trap (thanks to a bit of engineering work from me to fix the trap and a bit of bait preparation from HSO Koi) and had been involuntary signed up for today’s EVA to Candor Chasma for an empathetic release.

EVA #20 started with the Millenial Crew (read: Elena, Koi, and myself) returning a rock sample at Robert’s Rock Garden before taking our murine crew member to the Chasma. Our send-off was quite short… perhaps he had more important places to be? Fortunately, that was the only crew member we would lose on this EVA through the canyons on one of the most picturesque hikes we had performed on an EVA.

The rest of the crew was busy in the hab in the meantime. Crew Artist Gagnon was signing the back of the 1000-piece puzzle – a tradition that we weren’t aware of until it was completed. He made sure to incorporate our Phoenix crew logo and the infamous Monolith into a sketch next to the signature. Commander Laude virtually met with tomorrow’s CNBC representatives to discuss interview details.

The team reunited around noon and completed our final 100cameras module and go over logistics for tomorrow. Tomorrow’s segment is meant to address the question “Are We Ready for Mars?” Preparing for the answers to that question always sparks an enthralling conversation that has Crew Journalist Saavedra-Buckley furious jotting down notes for her article in shorthand.

The rest of our afternoon was spent watching the sequel to 2001: “2010: The Year We Made Contact”. 2 hours later, all of my Space Odyssey questions were answered, and I will leave this station more informed on historical sci-fi cinema. As the comms window opened, I went to repair our rear exterior airlock door, which has started to droop over the last week. Some quick hinge fastener tightening did the trick.

Speaking of engineering work, I wanted to end this final engineering-focused sol summary to talk about some of the conclusions we have drawn from two important pieces of our EVA equipment: our rovers and our space suits.

Regarding our rovers, we have four in our fleet: Curiosity, Spirit, Opportunity, and Perseverance. Spirit had battery issues and was out of commission for the duration of the mission, and Perseverance needed some parking brake and ignition repairs. We put a total of 22.8 hours of drive time among the Perseverance, Curiosity, and Opportunity rovers. And we preferred those vehicles in that same order, based on how efficiently their batteries’ life was consumed. Perseverance would use an average of 29% of its battery life per hour of use, Curiosity 36%, and Opportunity 50%. As I gathered this information, we used it to determine which rover was best for long range missions and also when we could use a less efficient battery to give Perseverance time to charge up.

In terms of EVA space suits, we have three flavors: two-piece suits (#1-6), one-piece suits (#7-10) and one experimental suit (#11). They all have pros and cons, and I’d like to summarize here when each should be used. The two-piece suits are the most realistic looking, provide the strongest airflow to the helmet, and are the most compact to easily ingress and egress from the rovers. However, it is the heaviest and most uncomfortable, tending to push your neck and head forward. The one-piece suit is the lightest and most comfortable, but it is the bulkiest and provides little ability to look around while driving. The experimental suit has an intermediate amount of weight and comfort and has the best helmet visibility, but it doesn’t have a battery gauge and can be uncomfortable to ride in a rover with (due to lack of neck support). In conclusion, I would recommend the two-piece suit for rover-intensive missions, the one-piece suit for walking-heavy missions, and the experimental suit for missions where you need to use a screen or closely inspect objects.

I hope this information is useful to future crews and Mission Support! We are off to get one last night of sleep in the habitat under simulation conditions. Tomorrow’s EVA #21 will be a culmination of our various projects and an opportunity to show others what we came here to do. Good night and ad astra!

Look Ahead Plan: Tomorrow we will be welcoming CNBC to the habitat for a news segment and taking them on an EVA to demonstrate our drone flights and 3D scanning capabilities. We will also begin shutting down the Science Dome and RAM and will consider ending simulation on Friday evening to get dinner in Hanksville.

Anomalies in work: None

Weather: Clear skies all day, temperatures ranging from 55 F to 65 F. Winds were light up to 5 mph, gusting up to 19 mph in the evening.

Crew Physical Status: All Healthy

EVA: Our 20th EVA was performed today, which consisted of the following objectives:

· Returning collected rock samples to Robert’s Rock Garden

· Releasing the captured mouse into the Martian landscape

· Exploring Candor Chasma

· Troubleshooting drone operations in the field

Reports to be filed: 1 EVA Request, 1 EVA Report, GreenHab Report, Operations Report, Journalist Report, Crew Photos

Support Requested: None

GreenHab Report – May 1st

Crew 315 GreenHab Report 01-05-2025

GreenHab Officer: Elena Saavedra Buckley

Environmental control (fan & heater): On

Average temperatures (last 24h): 83

Maximum temperature (last 24h): 95

Minimum temperature (last 24h): 71

Hours of supplemental light: Disabled

Daily water usage for crops: 4.2 gal

Daily water usage for research and/or other purposes: Misting mushrooms

Water in Blue Tank (200 gallon capacity): 90.1 gallons

Time(s) of watering for crops: 8:30 am

Changes to crops: None

Narrative: We’ve left most of the tomato plants, the herbs, and the beds near the water tank intact to give CNBC something green to film. After they’re done with that interview, we’ll finish tearing it down and leave a few tomato plants and herbs for Mission Support.

Harvest: None

Support/supplies needed: Do you want the herbs picked or left with roots so you can replant (or would you like us to plant them for you?)?

Operations Report – May 1st

Crew 315 Operations Report 01-05-2025

SOL: 11

Name of person filing report: Michael Andrews

Non-nominal systems: Catalytic Heater

Notes on non-nominal systems:

· (Bonus project) The catalytic heater near the exercise bike in the habitat does not work. It’s an engineering project to see if this can get properly started during this mission. Barring any more impactful repairs needed, this project will be worked tomorrow.

ROVERS

Spirit rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Opportunity rover used: Yes

Hours: 222.0 hours (0.8 hours used)

Beginning charge: 100%

Ending charge: 40%

Currently charging: No

Curiosity rover used: Yes

Hours: 318.3 hours (0.5 hours used)

Beginning charge: 100%

Ending charge: 67%

Currently charging: Yes

Perseverance rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

General notes on rovers: None

Summary of Hab operations: Team performed 1 EVA, completed the final 100cameras module, watched 2020: The Year We Made Contact, and prepared for tomorrow’s CNBC interview.

Water Use

Time of measurements: 1910
1) Per formula: 29.41 gallons
2) Smart Home Dashboard: 32.0 gallons

Water (static tank, remaining gallons): 170.97 gallons

Static tank pipe heater: Off

Static tank heater: Off

Toilet tank emptied: Yes

Summary of internet: Used to complete reports, learn languages, share team photos via Google drive, researching projects, and complete professional work.

Summary of suits and radios:

· EVA 20: Suits 9, 10, 11 and Radios 3, 4, 5

Summary of GreenHab operations: Watered plants and Crew Artist began the teardown GreenHab process by placing plants in trash bags.

WATER USE: 4.2 gallons

Heater: On

Supplemental light: Off

Harvest (name, weight in grams): None

Summary of ScienceDome operations: used for Commander’s meeting with CNBC, power source for Starlink Mini, and are for the Crew Artist to sign the back of the 1000-piece puzzle.

Dual split: Off

Summary of RAM operations: No tools used

Summary of any observatory issues: None

Summary of health and safety issues: None

Questions, concerns and requests to Mission Support: None

EVA Report – May 1st

Crew 315 EVA Report 01-05-2025

EVA # 20

Author: Michael Andrews

Purpose of EVA: To return rock samples collected at Robert’s Rock Garden for the 3D scanning project and further explore Candor Chasma based on results from EVA 19. At Candor Chasma we searched for more rock samples and performed more drone flight operations.

Start time: 0905

End time: 1150

Narrative: The crew departed the airlock with caught mouse in tow and headed to Robert’s Rock Garden. The last collected rock sample was quickly deposited in the location that it was collected from, and the team headed towards Candor Chasma.

At our destination, HSO Koi released the mouse (affectionately named Remi) in the surrounding environment. He quickly scurried out of sight and we began our hike through the Chasma.

The environment was technical at first to get to the riverbed below, but then quickly became flat and easy to traverse. We traveled down the chasma for about 35 minutes before stopping and heading back. No interesting rock samples were collected, and HSO Koi was not able to get the drone adequate signal for launch.

On the return trip, we stopped at a location above Compass Rock to survey it before exploring it for EVA #21 tomorrow. The team returned mouseless to the airlock at 1150.

Destination: Robert’s Rock Garden and Candor Chasma

Coordinates:

· Robert’s Rock Garden: Easting 518500, Northing 4249500

· Candor Chasma: Easting 520500, Northing 4251500

EVA Participants: Michael Andrews (EVA Lead and Crew Engineer), Urban Koi (HSO), and Elena Saavedra-Buckely (Crew Journalist)

Road(s) and routes per MDRS Map: Take Cow Dung Road south until reaching Robert’s Rock Garden, then complete first objective. Next, proceed north along Cow Dung Road until Galileo Road, then turn right. Proceed along Galileo Road until reaching Candor Chasma and setting out on foot (due east). To return to MDRS, take Galileo Road north and then Cow Dung Road south.

Mode of travel: Take one rover to the destinations above, then travel no more than 2 kilometers from the rovers on foot at each destination.

Vehicles you will be using: Opportunity and Curiosity

Supplemental Operations Report -May 4th

Date: 5/4/2025
Name of person filing report: David Steinhour
Reason for Report: Routine
Non-Nominal Systems: Crew car. Power system battery, inverters, generator. Robotic observatory. HAB outer shell. Spirit. Guest Wifi not turning on.

Power system: "Solar: The battery bank does not hold charge when sun is down and low on the horizon. Inverter Slave 1 and 2 went offline and do not restart, which limits us to 4kW when on solar.
Main generator has been monitored for oil leaks; minor leaks observed, Moreover, generator is consuming oil because of worn piston rings. Adding oil every day is necessary. Generator is limited to 8kW, see previous reports for details. No further leaks have been observed.
1) Oil, oil filter changed on 05/02/2025. Hours: 9484.3
2) Current hours – 9496.6"

Propane Readings: "Refilled 4/24
Station Tank: 79%
Director Tank: 75%
Intern Tank: 85%
Generator Tank: 53%"

Water: "Hab Static Tank – 170 gallons
GreenHab – 0 gallons
Outpost tank – 550 gallons"

Rovers: "Sojourner rover used: no
Hours: 213.4
Beginning Charge: 100 %
Ending Charge: 100 %
Currently Charging: No
Notes on Rovers: Spirit is not charging"

Cars: "Hab Car used and why, where: To Hanksville for supplies.
Crew Car used and why, where: Transporting crew to Grand Junction.
General notes and comments: Crew car driver’s side front ball joint is bad. Low oil pressure is most likely caused by a bad sensor (it does not change due to RPM or temperature like it should)."

Summary of Internet: Guest WiFi is not turning on despite all attempts, and most likely needs a firmware reset.
EVA suits and radios: "Suits: Nominal.
Comms: Headset #3 push-to-talk button nonfunctional.
T-Echo EVA-link: Nominal"
Campus wide inspection, if action taken, what and why: Nominal
Summary of Hab Operations: Small cracks in the wall of the Hab in the loft area are allowing cold air intrusion. Static tank and pipe heaters unplugged. Thermostat setup for 60F due to warm weather.
Summary of GreenHab Operations: Greenhab plants removed for summer temps. New potting soil resupplied.
Summary of SciDome Operations: Dual split is being turned on during day time for cooling.
Summary of Observatories Operations: Robotic observatory offline.
Summary of RAM Operations: Nominal
Summary of Outpost Operations: Nominal
Summary of Health and Safety Issues: Nominal

Copyright © The Mars Society. All rights reserved. | Main Site